OpenStudy (anonymous):

point T bisects UV. Find UV if UT=2 7/8

OpenStudy (anonymous):

bisect- dividing into 2 equal parts

OpenStudy (ghazi):

\[2\frac{ 7 }{ 8 }= (2*8)+7/8= 23/8\] now divide this by 2

OpenStudy (anonymous):

multiply not divide @ghazi

OpenStudy (ghazi):

yea...that was the mistake :(

OpenStudy (anonymous):

so I multiply 23/8 by 2?

OpenStudy (ghazi):

yes

OpenStudy (anonymous):



OpenStudy (anonymous):

23/4?

OpenStudy (ghazi):

yes

OpenStudy (anonymous):

thank you (: